# A Lie algebraic pattern behind logarithmic CFTs

### v2

@@ 2.4.4. Ramond sector of case (2)
$\begin{split}L_{0}&=\operatorname{Res}_{z}z^{2}Y(\Delta(\tfrac{\varpi_{r}}{\sqrt{2m-1}},z)\omega_{\mathrm{sh}}+e^{\Delta_{x}}\omega^{(s)},z)\\ &=\operatorname{Res}_{z}z^{2}Y(\omega_{\mathrm{sh}}+\omega^{(s)},z)+\tfrac{1}{16}+\operatorname{Res}_{z}z^{2}Y((\tfrac{(\varpi_{r})_{(1)}}{\sqrt{2m-1}}+\tfrac{1}{2}(\tfrac{(\varpi_{r})_{(1)}(\varpi_{r})_{(1)}-\sqrt{2m-1}(\varpi_{r})_{(2)}}{2m-1}))\omega_{\mathrm{sh}},z)\\ &=:\operatorname{Res}_{z}z^{2}Y(\omega_{\mathrm{sh}}+\omega^{(s)},z)+\tfrac{1}{16}+A(\alpha_{r})_{(0)}+B(\alpha_{r-1})_{(0)}+C,\end{split}$ (2.23)
where $\Delta_{x}$ is defined in [AM208], and $A$, $B$, $C$ are some constants. In the Ramond sector $\mathcal{S}_{R}V_{\lambda}$, we have
$\Delta_{\mu}=\tfrac{1}{2}|\mu|^{2}-\sqrt{2m-1}(1-\tfrac{1}{2m-1})(\mu,\rho)+A(\alpha_{r},\mu)+B(\alpha_{r-1},\mu)+C+\tfrac{1}{16}$ (2.24)
@@
    If $\mathfrak{g}=B_{r}$ and $p=2m-1$ for $m\in\mathbb{Z}_{\geq 1}$, then we have a shift system $(\Lambda,\uparrow,\{V_{\lambda}\}_{\lambda\in\Lambda})$, where
@@
Finally, the following is proved in the same manner as [Sug22, Lemma 2.18, 2.27] so we omit the proof.
@@
In Theorem 2.13(1) (resp. (2)), $V_{\lambda}^{\ast}(-2\rho)$ is isomorphic to $V_{w_{0}\ast\lambda^{\prime}}(-\rho^{\vee})$ (resp. $V_{w_{0}\ast\lambda^{\prime}}(-\rho)$) as $B$- and $V_{0}$-modules, where $\lambda^{\prime}$ is the representative of $-w_{0}(\lambda)$. In particular, for $0\leq n\leq l(w_{0})$, we have
@@ 3. Main results I
In this section, by applying Theorem 1.1 to the shift systems in Theorem 2.13, we show that the main results of [Sug21] and part of [Sug22] also hold for our cases. After stating the definitions and properties common to the two cases of Theorem 2.13, we will examine each case in Section 3.1 and 3.2, respectively.
@@
The multiplet $W$-(super)algebra is defined by the vertex operator (super) subalgebra
@@
On the other hand, in the same manner as [Sug22, Corollary 2.21], the 00-th sheaf cohomology
@@
of the sheaf associated with the $G$-equivariant vector bundle $G\times_{B}V_{0}$ over the flag variety $G/B$ has the VO(S)A structure induced from $V_{0}$, and the $n$-th sheaf cohomology $H^{n}(G\times_{B}V_{\lambda}(\mu))$ is an $H^{0}(G\times_{B}V_{0})$-module (for more detail, see [Sug22, Section 2.1] and the discussion just before [Sug22, Corollary 2.21]). Denote
$H^{0}(G\times_{B}V_{\lambda})\simeq\bigoplus_{\alpha\in P_{+}\cap Q}L_{\alpha+\lambda^{\bullet}}\otimes\mathcal{W}_{-\alpha+\lambda}$ (3.1)
@@
Let us use the same notation and setup in Theorem 2.13. The proofs are basically the same as in [Sug21, Sug22], so we will only outline them. By Theorem 1.1 and Theorem 2.13 we obtain the following.
@@
    (Feigin-Tipunin conjecture/construction) The evaluation map
@@
    (“Borel-Weil-Bott theorem”) For $\lambda\in\Lambda$ such that $(p\lambda_{\bullet}+\rho^{\vee},{}^{L}\theta)\leq p$ (resp. $(p\lambda_{\bullet}+\rho,{}^{L}\theta)\leq p$), we have a natural $G$- and $W_{\sqrt{p}Q}$-module isomorphism
    $H^{n}(G\times_{B}V_{\lambda})\simeq H^{n+l(w_{0})}(G\times_{B}V_{w_{0}\ast\lambda}(-\rho^{\vee}))\quad(\text{resp.}\quad H^{n}(G\times_{B}V_{\lambda})\simeq H^{n+l(w_{0})}(G\times_{B}V_{w_{0}\ast\lambda}(-\rho))).$
@@ 3.1. The non-super case (1)
[Sug21, Theorem 1.2, Lemma 4.21] For $\alpha\in P_{+}\cap Q$ and $\lambda\in\Lambda$ such that $(p\lambda_{\bullet}+\rho^{\vee},{}^{L}\theta)\leq p$,
$\operatorname{ch}\mathcal{W}_{-\alpha+\lambda}=\sum_{\sigma\in W}(-1)^{l(\sigma)}\operatorname{ch}V_{\sigma\ast\lambda}^{h=\alpha+\lambda^{\bullet}-\sigma\uparrow\lambda}=\sum_{\sigma\in W}(-1)^{l(\sigma)}\tfrac{q^{\frac{1}{2p}|-p\sigma(\alpha+\lambda^{\bullet}+\rho)+p\lambda_{\bullet}+\rho^{\vee}|^{2}}}{\eta(q)^{r}}.$ (3.2)
@@
[Sug22, Section 3.2] For $m\geq{}^{L}h^{\vee}-1$ and $\lambda=\lambda^{\bullet}$, $W_{\sqrt{p}Q}$ and $W_{\sqrt{p}P}$ are simple as VOA and generalized VOA, respectively, and $W_{\sqrt{p}(Q-\lambda^{\bullet})}$ is simple as $W_{\sqrt{p}Q}$-module. Furthermore, for $m\geq{}^{L}h^{\vee}-1$ and $\beta\in P_{+}$, $\mathcal{W}_{-\beta}$ is simple as $\mathcal{W}_{0}$-module.
@@ 3.2. The super case (2)
$\operatorname{ch}\mathcal{W_{-\alpha+\lambda}}=\operatorname{ch}F\sum_{\sigma\in W}(-1)^{l(\sigma)}\operatorname{ch}V_{\sigma\ast\lambda}^{h=\alpha+\lambda^{\bullet}-\sigma\uparrow\lambda}=\operatorname{ch}F\sum_{\sigma\in W}(-1)^{l(\sigma)}\tfrac{q^{\Delta_{\sqrt{p}(-(\alpha+\lambda^{\bullet})+\sigma\ast\lambda_{\bullet})}}}{\eta(q)^{r}},$ (3.3)
$\mathrm{sch}\mathcal{W}_{-\alpha+\lambda}=\mathrm{sch}F\sum_{\sigma\in W}(-1)^{l(\sigma)}\mathrm{sch}V_{\sigma\ast\lambda}^{h=\alpha+\lambda^{\bullet}-\sigma\uparrow\lambda}=\mathrm{sch}F\sum_{\sigma\in W}(-1)^{l(\sigma)+f(\sigma\circ(\alpha+\lambda^{\bullet}))}\tfrac{q^{\Delta_{\sqrt{p}(-(\alpha+\lambda^{\bullet})+\sigma\ast\lambda_{\bullet})}}}{\eta(q)^{r}},$ (3.4)
$\operatorname{ch}\mathcal{S}_{R}\mathcal{W}_{-\alpha+\lambda}=q^{C+\tfrac{1}{16}}\operatorname{ch}\iota^{*}F\sum_{\sigma\in W}(-1)^{l(\sigma)}\tfrac{q^{\Delta_{\sqrt{p}(-(\alpha+\lambda^{\bullet})+\sigma\ast\lambda_{\bullet})}+(A\alpha_{r}+B\alpha_{r-1},-\sqrt{p}(\sigma\circ(\alpha+\lambda^{\bullet})-\lambda_{\bullet})}}{\eta(q)^{r}},$ (3.5)
@@
For simply-laced $\mathfrak{g}$ (resp. rank $1$ case), modularity of the characters were already studied in [BrM17] (resp. [AM09, AM208]). We expect similar results hold in our cases too and plan to discuss in a future work.
@@ 4. Main results II
In Section 4, first we show that $\mathcal{W}_{0}$ is isomorphic to the corresponding principal $W$-algebra $\mathbf{W}^{k}(\mathfrak{g})$. It enables us to use another type character formula, i.e. Kazhdan-Lusztig type character formula (see [KT00, Theorem 1.1], [Ar07, Theorem 7.7.1]) and the discussion in Section 2.2 (see the last two paragraphs). Using these results, we will attempt to extend the simplicity theorems (Corollary 3.4 and Corollary 3.6) to $\lambda\in\Lambda$ satisfying condition (strong) (or, equivalently, the conditions in Lemma 2.9). In the non-super case (1), the simplicity theorem will be proved in the same manner as [Sug22]. On the other hand, the super case (2) has not been studied as much as the case (1), and we will prove the simplicity theorem under a particular assumption (exactness of $+$-reduction). Since the flow of the discussion is the same as that of [Sug22, Section 3.3], details are sometimes omitted.
@@ 4.1. Preliminary from W-algebra
Let us recall the notation in [Kac90, Chapter 6] (see also [Wak01, Chapter 1]) and [Ar07, ArF19, ACL19]. For an affine Lie algebra $\hat{\mathfrak{g}}$, the Cartan subalgebra $\hat{\mathfrak{h}}$ and its dual $\hat{\mathfrak{h}}^{\ast}$ are decomposed as $\hat{\mathfrak{h}}=\mathfrak{h}\oplus(\mathbb{C}K+\mathbb{C}d)$ and $\hat{\mathfrak{h}}^{\ast}=\mathfrak{h}^{\ast}\oplus(\mathbb{C}\delta+\mathbb{C}\Lambda_{0})$, respectively. For $\hat{\mu}\in\hat{\mathfrak{h}}^{\ast}$, we have
@@
By [Kac90, Proposition 6.5], the affine Weyl group $\hat{W}$ is $W\ltimes Q^{\vee}$ (for $\hat{\mathfrak{g}}=X^{(1)}_{r}$) and $W\ltimes Q_{B_{r}}$ (for $\hat{\mathfrak{g}}=A^{(2)}_{2r}$), respectively. The $\hat{W}$-action on $\hat{\mathfrak{h}}^{\ast}$ is given by
@@
For $k\in\mathbb{C}$, denote $\hat{\mathfrak{h}}^{\ast}_{k}$ the set of $\hat{\mu}\in\hat{\mathfrak{h}}^{\ast}$ such that $(\hat{\mu},K)=k$. Let us define $V^{k}(\mathfrak{g})=U(\hat{\mathfrak{g}})\otimes_{\mathfrak{g}[t]\oplus\mathbb{C}K}\mathbb{C}_{k}$ the universal affine vertex algebra at level $k$, where $\mathfrak{g}[t]$ (resp. $K$) act on $\mathbb{C}_{k}$ trivially (resp. $k\operatorname{id}$). More generally, we can define the Weyl module $\hat{V}_{\beta,k}=U(\hat{\mathfrak{g}})\otimes_{\mathfrak{g}[t]\oplus\mathbb{C}K}L_{\beta}$ over $V^{k}(\mathfrak{g})$ in the same manner. Clearly, we have $\hat{V}_{0,k}=V^{k}(\mathfrak{g})$. For $\hat{\mu}\in\hat{\mathfrak{h}}^{\ast}_{k}$, define $\hat{M}(\hat{\mu})=U(\hat{\mathfrak{g}})\otimes_{U(\hat{\mathfrak{b}})}\mathbb{C}_{\hat{\mu}}$ the Verma module and its simple quotient $\hat{L}(\hat{\mu})$, respectively, where $U(\hat{\mathfrak{n}})$ (resp. $\mathfrak{h}$, $K$) act on $\mathbb{C}_{\hat{\mu}}$ trivially (resp. by the character of $\mu$, $k\operatorname{id}$).
By applying the Drinfeld-Sokolov reduction $H^{0}_{\mathrm{DS}}(-)$ to $V^{k}(\mathfrak{g})$, the (universal) principal $W$-algebra $\mathbf{W}^{k}(\mathfrak{g})$ at level $k$ and its unique simple quotient $\mathbf{W}_{k}(\mathfrak{g})$ are obtained. Note that we have the Feigin-Frenkel duality
$\mathbf{W}^{k}(\mathfrak{g})\simeq\mathbf{W}^{{}^{L}k}({}^{L}\mathfrak{g}),\ \ r^{\vee}(k+h^{\vee})({}^{L}k+{}^{L}h^{\vee})=1,\ \ \mathfrak{g}=X_{r}^{(1)},$ (4.1)
@@
in our cases (see [FF91a, FF91b], [Gen17, Remark 6.5]). More generally, for $\check{\mu}\in\check{P}$ in the coweight lattice $\check{P}=Q^{\ast}$, we have the twisted Drinfeld-Sokolov reduction $H^{0}_{\mathrm{DS},\check{\mu}}(-)$ and the Arakawa-Frenkel module
@@
over $\mathbf{W}^{k}(\mathfrak{g})$. For the Langlands dual ${}^{L}\mathfrak{g}$, we can define $\check{\mathbf{T}}^{k}_{\check{\mu},\beta}=H^{0}_{\mathrm{DS},\beta}({}^{L}\hat{V}_{\check{\mu},k})$ in the same manner. Then for $k\not\in Q$ or the case in [Sug22], the Feigin-Frenkel duality $\mathbf{T}^{k+h^{\vee}}_{\beta,\check{\mu}}\simeq\check{\mathbf{T}}^{{}^{L}k+{}^{L}h^{\vee}}_{\check{\mu},\beta}$ under the identification $\mathbf{W}^{k}(\mathfrak{g})\simeq\mathbf{W}^{{}^{L}k}({}^{L}\mathfrak{g})$ is known to hold (see [ArF19, Sug22]). Let $\chi_{\mu}:Z(\mathfrak{g})\rightarrow\mathbb{C}$ with $\mu\in\mathfrak{h}^{\ast}$ be the map in [ACL19, (27)], where $Z(\mathfrak{g})$ is the center of the universal enveloping algebra of $U(\mathfrak{g})$. Then one can define the Verma module $\mathbf{M}_{k}(\chi_{\mu}):=U(\mathbf{W}^{k}(\mathfrak{g}))\otimes_{U(\mathbf{W}^{k}(\mathfrak{g})_{\geq 0})}\mathbb{C}_{\chi_{\mu}}$ over $\mathbf{W}^{k}(\mathfrak{g})$ with highest weight $\chi_{\mu}$ and its simple quotient $\mathbf{L}_{k}(\chi_{\mu})$ in the same manner above (in the Langlands dual case, denote $\check{\mathbf{M}}_{k}(\chi_{\mu})$ and $\check{\mathbf{L}}_{k}(\chi_{\mu})$, respectively).

### v3

@@ 2.4.4. Ramond sector of case (2)
(2.23) $\begin{split}L_{0}&=\operatorname{Res}_{z}z^{2}Y(\Delta(\tfrac{\varpi_{r}}{\sqrt{2m-1}},z)\omega_{\mathrm{sh}}+e^{\Delta_{x}}\omega^{(s)},z)\\ &=\operatorname{Res}_{z}z^{2}Y(\omega_{\mathrm{sh}}+\omega^{(s)},z)+\tfrac{1}{16}+\operatorname{Res}_{z}z^{2}Y((\tfrac{(\varpi_{r})_{(1)}}{\sqrt{2m-1}}+\tfrac{1}{2}(\tfrac{(\varpi_{r})_{(1)}(\varpi_{r})_{(1)}-\sqrt{2m-1}(\varpi_{r})_{(2)}}{2m-1}))\omega_{\mathrm{sh}},z)\\ &=:\operatorname{Res}_{z}z^{2}Y(\omega_{\mathrm{sh}}+\omega^{(s)},z)+\tfrac{1}{16}+A(\alpha_{r})_{(0)}+B(\alpha_{r-1})_{(0)}+C,\end{split}$
where $\Delta_{x}$ is defined in [8], and $A$, $B$, $C$ are some constants. In the Ramond sector $\mathcal{S}_{R}V_{\lambda}$, we have
(2.24) $\Delta_{\mu}=\tfrac{1}{2}|\mu|^{2}-\sqrt{2m-1}(1-\tfrac{1}{2m-1})(\mu,\rho)+A(\alpha_{r},\mu)+B(\alpha_{r-1},\mu)+C+\tfrac{1}{16}$
@@
    If $\mathfrak{g}=B_{r}$ and p=2m1p=2m-1 for $m\in\mathbb{Z}_{\geq 1}$, then we have a shift system $(\Lambda,\uparrow,\{V_{\lambda}\}_{\lambda\in\Lambda})$, where
@@
Finally, the following is proved in the same manner as [46, Lemma 2.18, 2.27] so we omit the proof.
@@
In Theorem 2.13(1) (resp. (2)), $V_{\lambda}^{\ast}(-2\rho)$ is isomorphic to $V_{w_{0}\ast\lambda^{\prime}}(-\rho^{\vee})$ (resp. $V_{w_{0}\ast\lambda^{\prime}}(-\rho)$) as $B$- and $V_{0}$-modules, where $\lambda^{\prime}$ is the representative of $w_{0}(\lambda)$. In particular, for $0\leq n\leq l(w_{0})$, we have
@@ 3. Main results I
In this section, by applying Theorem 1.1 to the shift systems in Theorem 2.13, we show that the main results of [45] and part of [46] also hold for our cases. After stating the definitions and properties common to the two cases of Theorem 2.13, we will examine each case in Section 3.1 and 3.2, respectively.
@@
The multiplet WW-(super)algebra is defined by the vertex operator (super) subalgebra
@@
On the other hand, in the same manner as [46, Corollary 2.21], the 0-th sheaf cohomology
@@
of the sheaf associated with the $G$-equivariant vector bundle $G\times_{B}V_{0}$ over the flag variety $G/B$ has the VO(S)A structure induced from $V_{0}$, and the $n$-th sheaf cohomology $H^{n}(G\times_{B}V_{\lambda}(\mu))$ is an $H^{0}(G\times_{B}V_{0})$-module (for more detail, see [46, Section 2.1] and the discussion just before [46, Corollary 2.21]). Denote
(3.1) $H^{0}(G\times_{B}V_{\lambda})\simeq\bigoplus_{\alpha\in P_{+}\cap Q}L_{\alpha+\lambda^{\bullet}}\otimes\mathcal{W}_{-\alpha+\lambda}$
@@
Let us use the same notation and setup in Theorem 2.13. The proofs are basically the same as in [45, 46], so we will only outline them. By Theorem 1.1 and Theorem 2.13 we obtain the following.
@@
    The evaluation map
@@
    For $\lambda\in\Lambda$ such that $(p\lambda_{\bullet}+\rho^{\vee},{}^{L}\theta)\leq p$ (resp. $(p\lambda_{\bullet}+\rho,{}^{L}\theta)\leq p$), we have a natural $G$- and $W_{\sqrt{p}Q}$-module isomorphism
    $H^{n}(G\times_{B}V_{\lambda})\simeq H^{n+l(w_{0})}(G\times_{B}V_{w_{0}\ast\lambda}(-\rho^{\vee}))$
    $(\text{resp.}\quad H^{n}(G\times_{B}V_{\lambda})\simeq H^{n+l(w_{0})}(G\times_{B}V_{w_{0}\ast\lambda}(-\rho))).$
@@ 3.1. The non-super case (1)
[45, Theorem 1.2, Lemma 4.21] For $\alpha\in P_{+}\cap Q$ and $\lambda\in\Lambda$ such that $(p\lambda_{\bullet}+\rho^{\vee},{}^{L}\theta)\leq p$,
(3.2) $\operatorname{ch}\mathcal{W}_{-\alpha+\lambda}=\sum_{\sigma\in W}(-1)^{l(\sigma)}\operatorname{ch}V_{\sigma\ast\lambda}^{h=\alpha+\lambda^{\bullet}-\sigma\uparrow\lambda}=\sum_{\sigma\in W}(-1)^{l(\sigma)}\tfrac{q^{\frac{1}{2p}|-p\sigma(\alpha+\lambda^{\bullet}+\rho)+p\lambda_{\bullet}+\rho^{\vee}|^{2}}}{\eta(q)^{r}}.$
@@
[46, Section 3.2] For $m\geq{}^{L}h^{\vee}-1$ and $\lambda=\lambda^{\bullet}$, $W_{\sqrt{p}Q}$ and $W_{\sqrt{p}P}$ are simple as VOA and generalized VOA, respectively, and $W_{\sqrt{p}(Q-\lambda^{\bullet})}$ is simple as $W_{\sqrt{p}Q}$-module. Furthermore, for $m\geq{}^{L}h^{\vee}-1$ and $\beta\in P_{+}$, $\mathcal{W}_{-\beta}$ is simple as $\mathcal{W}_{0}$-module.
@@ 3.2. The super case (2)
(3.3) $\operatorname{ch}\mathcal{W_{-\alpha+\lambda}}=\operatorname{ch}F\sum_{\sigma\in W}(-1)^{l(\sigma)}\operatorname{ch}V_{\sigma\ast\lambda}^{h=\alpha+\lambda^{\bullet}-\sigma\uparrow\lambda}=\operatorname{ch}F\sum_{\sigma\in W}(-1)^{l(\sigma)}\tfrac{q^{\Delta_{\sqrt{p}(-(\alpha+\lambda^{\bullet})+\sigma\ast\lambda_{\bullet})}}}{\eta(q)^{r}},$
(3.4) $\mathrm{sch}\mathcal{W}_{-\alpha+\lambda}=\mathrm{sch}F\sum_{\sigma\in W}(-1)^{l(\sigma)}\mathrm{sch}V_{\sigma\ast\lambda}^{h=\alpha+\lambda^{\bullet}-\sigma\uparrow\lambda}=\mathrm{sch}F\sum_{\sigma\in W}(-1)^{l(\sigma)+f(\sigma\circ(\alpha+\lambda^{\bullet}))}\tfrac{q^{\Delta_{\sqrt{p}(-(\alpha+\lambda^{\bullet})+\sigma\ast\lambda_{\bullet})}}}{\eta(q)^{r}},$
(3.5) $\operatorname{ch}\mathcal{S}_{R}\mathcal{W}_{-\alpha+\lambda}=q^{C+\tfrac{1}{16}}\operatorname{ch}\iota^{*}F\sum_{\sigma\in W}(-1)^{l(\sigma)}\tfrac{q^{\Delta_{\sqrt{p}(-(\alpha+\lambda^{\bullet})+\sigma\ast\lambda_{\bullet})}+(A\alpha_{r}+B\alpha_{r-1},-\sqrt{p}(\sigma\circ(\alpha+\lambda^{\bullet})-\lambda_{\bullet})}}{\eta(q)^{r}},$
@@
For simply-laced $\mathfrak{g}$ (resp. rank 11 case), modularity of the characters were already studied in [11] (resp. [9, 8]). We expect similar results hold in our cases too.
@@ 4. Main results II
In Section 4, first we show that $\mathcal{W}_{0}$ is isomorphic to the corresponding principal W-algebra $\mathbf{W}^{k}(\mathfrak{g})$. It enables us to use another type character formula, i.e. Kazhdan–Lusztig type character formula (see [37, Theorem 1.1], [10, Theorem 7.7.1]) and the discussion in Section 2.2 (see the last two paragraphs). Using these results, we will attempt to extend the simplicity theorems (Corollary 3.4 and Corollary 3.6) to $\lambda\in\Lambda$ satisfying condition (strong) (or, equivalently, the conditions in Lemma 2.9). In the non-super case (1), the simplicity theorem will be proved in the same manner as [46]. On the other hand, the super case (2) has not been studied as much as the case (1), and we will prove the simplicity theorem under a particular assumption (exactness of ++-reduction). Since the flow of the discussion is the same as that of [46, Section 3.3], details are sometimes omitted.
@@ 4.1. Preliminary from W-algebra
Let us recall the notation in [34, Chapter 6] (see also [50, Chapter 1]) and [10, 5, 6]. For an affine Lie algebra $\hat{\mathfrak{g}}$, the Cartan subalgebra $\hat{\mathfrak{h}}$ and its dual $\hat{\mathfrak{h}}^{\ast}$ are decomposed as $\hat{\mathfrak{h}}=\mathfrak{h}\oplus(\mathbb{C}K+\mathbb{C}d)$ and $\hat{\mathfrak{h}}^{\ast}=\mathfrak{h}^{\ast}\oplus(\mathbb{C}\delta+\mathbb{C}\Lambda_{0})$, respectively. For $\hat{\mu}\in\hat{\mathfrak{h}}^{\ast}$, we have
@@
[…]
@@
By applying the Drinfeld–Sokolov reduction $H^{0}_{\mathrm{DS}}(-)$ to $V^{k}(\mathfrak{g})$, the (universal) principal W-algebra $\mathbf{W}^{k}(\mathfrak{g})$ at level $k$ and its unique simple quotient $\mathbf{W}_{k}(\mathfrak{g})$ are obtained. Note that we have the Feigin–Frenkel duality
(4.1) $\mathbf{W}^{k}(\mathfrak{g})\simeq\mathbf{W}^{{}^{L}k}({}^{L}\mathfrak{g}),\ \ r^{\vee}(k+h^{\vee})({}^{L}k+{}^{L}h^{\vee})=1,\ \ \mathfrak{g}=X_{r}^{(1)},$
(4.2) $\mathbf{W}^{k}(\mathfrak{osp}(1|2r))\simeq\mathbf{W}^{{}^{L}k}(\mathfrak{osp}(1|2r)),\ \ 4(k+h^{\vee})({}^{L}k+h^{\vee})=1$
in our cases (see [22, 23], [28, Remark 6.5]). More generally, for $\check{\mu}\in\check{P}$ in the coweight lattice $\check{P}=Q^{\ast}$, we have the twisted Drinfeld–Sokolov reduction $H^{0}_{\mathrm{DS},\check{\mu}}(-)$ and the Arakawa–Frenkel module
@@
over $\mathbf{W}^{k}(\mathfrak{g})$. For the Langlands dual $^{L}\mathfrak{g}$, we can define $\check{\mathbf{T}}^{{}^{L}k+{}^{L}h^{\vee}}_{\check{\mu},\beta}=H^{0}_{\mathrm{DS},\beta}(\hat{V}_{\beta,{}^{L}k})$ in the same manner. Then for $k\not\in Q$ or the case in [46], the Feigin–Frenkel duality $\mathbf{T}^{k+h^{\vee}}_{\beta,\check{\mu}}\simeq\check{\mathbf{T}}^{{}^{L}k+{}^{L}h^{\vee}}_{\check{\mu},\beta}$ under the identification $\mathbf{W}^{k}(\mathfrak{g})\simeq\mathbf{W}^{{}^{L}k}({}^{L}\mathfrak{g})$ is known to hold (see [5, 46]). Let $\chi_{\mu}:Z(\mathfrak{g})\rightarrow\mathbb{C}$ with $\mu\in\mathfrak{h}^{\ast}$ be the map in [6, (27)], where $Z(\mathfrak{g})$ is the center of the universal enveloping algebra of $U(\mathfrak{g})$. Then one can define the Verma module $\mathbf{M}_{k}(\chi_{\mu}):=U(\mathbf{W}^{k}(\mathfrak{g}))\otimes_{U(\mathbf{W}^{k}(\mathfrak{g})_{\geq 0})}\mathbb{C}_{\chi_{\mu}}$ over $\mathbf{W}^{k}(\mathfrak{g})$ with highest weight $\chi_{\mu}$ and its simple quotient $\mathbf{L}_{k}(\chi_{\mu})$ in the same manner above (in the Langlands dual case, denote $\check{\mathbf{M}}_{k}(\chi_{\mu})$ and $\check{\mathbf{L}}_{k}(\chi_{\mu})$, respectively).